OpenStudy (anonymous):

antiderivative of (1/x^2) dx?

zepdrix (zepdrix):

Rewrite it using negative exponents, then apply the power rule for integration! :) It shouldn't be too bad :D Lemme know if any of that is a little confusing.

OpenStudy (anonymous):

i have got x^-2 dx = x^-1/-1 + c, is that right?

zepdrix (zepdrix):

Yes looks good ! ^^

zepdrix (zepdrix):

Last step might be to rewrite it without the fraction part though, so it looks a little cleaner :)\[-x^{-1}+c\]

OpenStudy (anonymous):

i have 7 more antiderivative problems...I have attempted all of them I just need someone to verify. would you be willing to help me?

zepdrix (zepdrix):

sure, just post your next one in here if you like. I'll try to help if I have time :D But I'm sure someone can get to you if I don't! :D

OpenStudy (anonymous):

ok next one is similar to first antiderivative of (1/x) dx

zepdrix (zepdrix):

Unfortunately, the power rule for integration doesn't apply to x^-1. Remember what this special one gives you? :O

OpenStudy (anonymous):

what exactly is the power rule of integration btw

zepdrix (zepdrix):

\[\huge \int\limits_{}^{}x^n dx=\frac{ x^{n+1} }{ n+1 }+c\]

zepdrix (zepdrix):

You increase the exponent by one, then divide by the NEW power.

OpenStudy (anonymous):

how does this explain the -1 in the denominator of the first problem

zepdrix (zepdrix):

\[\huge \int\limits_{}^{}x^{-2}dx=\frac{ x^{-2+1} }{ -2+1 }+c=\frac{ x^{-1} }{ -1 }+c\]

OpenStudy (anonymous):

i see how now. thanks

OpenStudy (anonymous):

what about the second one, what would you do since you cant use power rule of integration

zepdrix (zepdrix):

It's one that you'll end up memorizing. If you don't know what it is, you need to go back to differentiation. \[\huge (\ln x)' = \frac{ 1 }{ x }\]\[\huge \int\limits_{}^{}\frac{ 1 }{ x }dx=?\]

OpenStudy (anonymous):

ln[x] + c?

zepdrix (zepdrix):

yes good c: Now there is something going on with the domain when you take this integral. So when you deal with DEFINITE integrals, you want to remember that it's actually producing \[\huge \ln |x| + c\] Absolute x.

zepdrix (zepdrix):

But that's not something we'll really need to worry about for indefinite integrals :D

OpenStudy (anonymous):

ok thanks for that one next one is antiderivative of x^3 (x^4 + 1) ^2 dx

zepdrix (zepdrix):

So this is where you get an introduction to U substitutions. Hmm

zepdrix (zepdrix):

See how we have an x^4, and an x^3? We can use the fact that they're one power apart to think of it as (Something)... and the derivative of (Something) Let's apply a u substitution. Let \[\huge u=x^4+1\] Taking the derivative of both sides with respect to x... \[\huge \frac{ du }{ dx }=3x^3\]

OpenStudy (anonymous):

wouldnt the derivative be 4x^3

zepdrix (zepdrix):

Yes thank you XD sorry about that.

OpenStudy (anonymous):

ok no problem

zepdrix (zepdrix):

From here, there is a process that allows you to rewrite the dx on the other side next to the 4x^3. Just simply think of it as multiplication. That's not entirely accurate, but for our purposes it is fine to think of it that way :D \[\huge du=4x^3dx\]

OpenStudy (anonymous):

for this part i wrote it as 1/4du = x^3 dx

zepdrix (zepdrix):

K good, you're thinking ahead c:

zepdrix (zepdrix):